# Trajectory Prediction for Autonomous Driving: Progress, Limitations, and Future Directions

### v2

@@ Abstract
As the potential for autonomous vehicles to be integrated on a large scale into modern traffic systems continues to grow, ensuring safe navigation in dynamic environments is crucial for smooth integration. To guarantee safety and prevent collisions, autonomous vehicles must be capable of accurately predicting the trajectories of surrounding traffic agents. Over the past decade, significant efforts from both academia and industry have been dedicated to designing solutions for precise trajectory forecasting. These efforts have produced a diverse range of approaches, raising questions about the differences between these methods and whether trajectory prediction challenges have been fully addressed. This paper reviews a substantial portion of recent trajectory prediction methods proposing a taxonomy to classify existing solutions. A general overview of the prediction pipeline is also provided, covering input and output modalities, modeling features, and prediction paradigms existing in the literature. In addition, the paper discusses active research areas within trajectory prediction, addresses the posed research questions, and highlights the remaining research gaps and challenges.
@@ 1 Introduction
The era of autonomous driving is on the verge of revolutionizing the modern transportation system. Automotive companies like Tesla, Waymo and Cruise are pioneers in today’s self-driving vehicle technology. The practical, innovative solutions designed by these companies, in conjunction with academic research, have led to rapid development in perception, prediction, planning, and control algorithms for autonomous driving. Predicting the future trajectories of traffic participants is crucial for ensuring safety in autonomous driving. The predicted trajectories are passed as input to planning algorithms to exclude paths leading to potential collisions and to select a safe course of action. Due to the diversity of traffic scenarios, the complex interactions between traffic participants, various maneuvers, and computational time requirements, the trajectory prediction problem poses a range of challenges (please refer to Figure 1 for factors impacting trajectory forecasting).
To accurately predict trajectories, the prediction model must account for the motion of vehicles, cyclists, pedestrians — all traffic participants. For vehicle motion, autonomous vehicles (AV) should be able to track surrounding vehicles and their potential maneuvers. As human drivers have an intuition of inter-dependency of surrounding vehicles, e.g., it can be anticipated if a neighboring car is cut off, there is a high chance that this vehicle will be compelled to perform a risky maneuver and potentially interfere with the ego-vehicle’s lane to avoid collision or suddenly hit the breaks. To equip autonomous vehicles with interaction-aware intuition, the prediction module needs to build a representation of each target agent’s surroundings and accounting for inter-dependency of traffic participants’ decisions. Regarding the human motion, beyond the complexity of interaction and inter-dependency, predicting human trajectories involves accounting for complex human behavior. Pedestrians’ motion is influenced not only by road topology and actions of surrounding agents but also by social rules, norms, and relations between agents. To accurately predict pedestrian trajectories, many factors need to be inferred from contextual and general knowledge information, e.g., a child and an adult exhibit different motion patterns. Additionally, prediction models that rely on traffic rules benefit from a reduced set of possible future trajectories. However, human drivers often violate traffic rules by over-speeding, changing lanes through a solid line, turning from a non-designated lane, ignoring stop signs, or even passing through a red light. Prediction models need to account for the risk of such violations to ensure robust and reliable predictions in real-world scenarios. Lastly, the trajectory prediction task has a multimodal nature, i.e., the same past trajectory can unfold into several possible future trajectories, e.g., a vehicle approaching an intersection in a lane marked for both straight and left-turn movements may continue forward or make a left turn. To account for multimodality, the output of the prediction model must contain the distribution of possible trajectories with probabilities.
@@
To address these challenges, the research community has focused on designing models capable of capturing complex patterns essential for accurate prediction. Among the top-performing solutions is Trajectron++ [1], which integrates environmental context through a graph-structured recurrent architecture and Spatio-Temporal Transformer Network (S2TNet) [2] leveraging transformer layers to capture both spatial and temporal dependencies. Further, from the intersection between academia and industry, the MultiPath++ model [3], released by the Waymo research group, employs multi-context gating fusion and dynamically learns reference points (anchors) to support goal-oriented predictions. Collectively, these models combine diverse techniques, such as graph-based interaction modeling, transformer architectures, and dynamic goal sampling, leading to more accurate and robust trajectory predictions. On the other side, the commercial research prioritizes real-time execution, cost-effectiveness, and handling uncertain situations. This leads each company to use its own combination of sensors, compatible with the set of algorithms it employs. Tesla’s prediction algorithm is an end-to-end vision-based deep learning model [^1] https://www.tesla.com/support/autopilot, forecasting trajectories based on the visual sequences and aggregated contextual information over time. The model continuously learns from the vast amount of data collected by Tesla’s global fleet of vehicles. In contrast, Waymo utilizes a combination of high-fidelity sensors (LiDAR, cameras, and radar). The sensor data is processed by the perception module and then passed to motion prediction and planning components. Waymo adopts a modular structure, ensuring that each component can be thoroughly tested independently and then integrated into a single pipeline [^2] https://waymo.com/research/. It is challenging to determine the specific algorithms employed in real operations, but Waymo has a strong R&D focus, actively publishing its latest research findings. Aurora company focuses on autonomous solutions for trucks serving logistics purposes. Its prediction module [^3] https://blog.aurora.tech/ performs joint object detection and trajectory forecasting, combining learning-based methods with rule-based constraints to align predictions with driving rules. To ensure safety across a wide range of scenarios, from highway driving delivering cargo between cities to urban logistics, Aurora’s system is tested extensively through a combination of virtual testing suites and real-world operations. Each company’s strategy reflects a balance between predictive accuracy and operational constraints, tailored to the specific purpose of their vehicles.
@@
On a testing side of autonomous driving algorithms, two robust open-source platforms provide environment for evaluation: Autoware [^4] https://github.com/autowarefoundation/autoware and Apollo [^5] https://github.com/ApolloAuto/apollo. Autoware’s perception stack includes functionalities such as object recognition (2D/3D detection using LiDAR, cameras, and radar), obstacle segmentation (distinguishing drivable areas via point cloud processing), occupancy grid mapping (representing occupied and free spaces), and traffic light recognition (detecting light states and directions). However, it lacks an integrated prediction module. Starting from release 5.0, Apollo incorporates an integrated prediction module [^6] https://github.com/ApolloAuto/apollo/blob/master/modules/prediction/README.md, supporting a range of predictors: Free Move Predictor for non-lane-constrained obstacles, Single Lane Predictor for vehicles on a single lane, and Lane Sequence Predictor for multi-lane and behavior-based trajectory predictions. Later versions, such as 6.0.0 and 7.0.0, introduced advanced prediction models, including a semantic map-based pedestrian prediction model and the Inter-TNT prediction model with interactive prediction and planning evaluation. While Apollo provides a comprehensive platform for testing prediction algorithms within the full autonomous driving stack, there is a clear need for an open, lightweight platform that enables faster experimentation with prediction modules. Such a platform should integrate a perception component to provide realistic inputs for testing, while remaining flexible enough to support subsequent integration into full-stack systems.
@@
To address the posed research questions, this study surveys the latest advancements in prediction models and offers a structured framework for understanding the trajectory forecasting landscape. We comprehensively review key factors that influence the performance and robustness of prediction models. This analysis aims to assist researchers in identifying gaps in the current literature and guiding future developments in the field. Our contributions are as follows:
@@
    Provide an open repository [^7] https://github.com/AV-Lab/Trajectory_Prediction_Models_Zoo featuring two prediction pipelines to facilitate experimentation, benchmarking, and the development of new methods. The first pipeline includes multiple learning-based models that can be trained on SOTA or custom datasets, enabling flexibility for diverse research needs. The second pipeline implements a detect–track–predict approach that processes inputs from various sensor configurations (e.g., frontal camera, stereo camera, LiDAR), enabling emulation of real-world deployment scenarios.
@@ 1.1 Related Surveys
Figure 2: Structure of the review paper: the study comprises seven main sections. The first two sections - an overview of the prediction pipeline and a survey of modeling approaches, provide general knowledge of the trajectory prediction domain. Next two sections are dedicated to incorporating driving knowledge and addressing uncertainty focusing enhancing robustness and generalization of prediction models. Last three sections cover planning-conditioned prediction, VLM-guided prediction, and collaboration between intelligent vehicles for trajectory prediction.
@@ 1.2 Scope and Structure
The scope of this review is defined by the objective to offer a comprehensive overview of the trajectory prediction field for novice researchers and already experienced practitioners. Figure 2 illustrates the structure of our review, detailing the topics and main points discussed.
The first part consists of two sections aimed at providing general knowledge of the trajectory prediction realm. Section 2 provides an overview of a trajectory prediction pipeline and covers problem formulation and prediction paradigms. Section 3 lays out the taxonomy and categorization of the modeling approaches and elaborates on the models available in the literature from each category. The next part is dedicated to the discussion of factors impacting model robustness and generalization. This part also comprises two sections. Section 4 reviews the inclusion of driving knowledge in prediction models, covering different types of driving knowledge and the ways they are incorporated into prediction models. Section 5 is dedicated to uncertainty estimation and mitigation in prediction models. It elaborates on the sources of uncertainty and approaches to mitigate the risks associated with them. The last part depicts active research areas and is presented by three sections. Section 6 discusses planning-prediction-coupled frameworks, where for generating predictions the ego-vehicle models internal planning intentions of surrounding agents. Next, Section 7 discusses the application of Vision-Language models in scene understanding and trajectory prediction. Further, Section 8 discusses the concept of collaborative perception, elaborating on different types of fusion and different types of challenges appearing because of collaborating systems. Lastly, Section 9 provides a discussion on the posed research questions, outlines the research gaps, and highlights future research directions.
@@ 2.1 Problem Formulation
In the given formulation, the predictor ($\mathcal{P}$) is a function of a single parameter, past trajectories ($\mathcal{T}$). This can be expanded to make $\mathcal{P}$ a function of multiple parameters if additional inputs are available, e.g., road geometrical features ($G$) or interaction modeling features ($I$).
@@ 2.2.1 Past Trajectories
Past trajectory is a fundamental input in trajectory prediction models. Predicting future trajectories solely based on past trajectories [17, 18, 19, 20] forces the models to learn patterns by relying purely on spatio-temporal dependencies. This approach, which excludes additional contextual or environmental information, is efficient in terms of prediction speed and computational cost. However, it lacks the capacity to capture knowledge about the environment or interactions between agents, making it less robust in complex scenarios. One way to implicitly encode interaction of the target agent is to include the past trajectories of its surrounding agents in the prediction process, as shown in the studies [21, 22, 23, 24, 25, 26, 27]. This enables predictions that are more reflective of real-world behaviors, such as reactions to nearby vehicles or pedestrians. However, the main limitation of these models is reliance on one input and assumption of error-free observations. The models are trained using ground-truth past trajectories, whereas in practical applications, the past trajectories available to prediction models are often noisy or incomplete, and robust handling of these inaccuracies is critical for practical deployment. Typically the agent state in past trajectories is represented as $(x,y)$ coordinates. Additionally, heading and speed can be added complementary. Studies [28, 29, 30] have highlighted the importance of including velocity, acceleration, and orientation as input features to enhance prediction accuracy.
@@ 2.2.2 Sensors Data
Sensor data, provided by exteroceptive sensors such as LiDAR, radar, and cameras, enables prediction models to capture environmental context for trajectory prediction models. This section first examines each sensor individually and then discusses approaches of integrating multi-modal inputs.
@@
LiDAR. LiDAR sensors enable environmental perception through laser-based distance measurements, making them indispensable for autonomous driving systems. Two types of LiDAR are commonly employed: perimeter LiDAR and $360^{\circ}$ LiDAR. Perimeter LiDAR, mounted on the vehicle’s edges or corners, is optimized for monitoring close-range objects, blind spots, and collision risks, particularly in scenarios like parking or navigating tight spaces. In contrast, $360^{\circ}$ LiDAR, typically roof-mounted, provides a complete $360^{\circ}$ field of view, supporting long-range detection and situational awareness. Current SOTA datasets predominantly include data from $360^{\circ}$ LiDARs (please refer to Figure 4 for examples), while perimeter LiDARs are used by automotive manufacturers as complementary sensors to enhance overall perception. One of LiDAR’s key strengths is its ability to function reliably in all lighting conditions, including complete darkness and bright sunlight, due to its light-invariant nature. However, its performance degrades in adverse weather, such as heavy rain, or snow, where laser beam scattering reduces detection range and accuracy along with highly reflective surfaces such as glass, potentially creating blind spots. Despite these challenges, LiDAR remains highly effective for detecting obstacles [31], vehicles [32, 33, 34], lane markers [35, 36, 37, 38], and forecasting [39].
@@
Two significant limitations hinder the widespread adoption of LiDAR technology: its relatively high cost and its accuracy in detecting pedestrians. Although LiDAR technology is advancing and prices are gradually declining, it remains expensive and impractical for consumer-level mass-market adoption. The second limitation is its lower accuracy in recognizing fine-grained classes such as pedestrians and cyclists. This disparity stems from its sparse, geometry-centric output and lack of rich semantic information, which makes interpretation of complex objects challenging. For example, on the KITTI leaderboard, the top pedestrian detection score is only 56% [42], compared to a 93% vehicle detection score achieved by [43]. As a result, LiDAR systems are ideally complemented by cameras, which provide the necessary semantic context to enhance pedestrian detection performance.
Camera. Cameras play a crucial role in perception of autonomous vehicles in detecting traffic participants, road signs, and traffic lights. These sensors are characterized by their resolution, field of view, and configuration. The vision community is highly advanced, providing efficient algorithms for object detection [44], semantic segmentation [45], and lane recognition [46]. Camera setups in AV systems are generally categorized as single frontal cameras [47] or multi-view configurations [48]. Single frontal camera setups primarily rely on 2D perception and are mostly used in driving assistance systems. In contrast, multi-view setups [49] arrange cameras to ensure complementary, non-overlapping fields of view, providing 360-degree observability. Additionally, camera systems employ either monocular or stereo setups. Monocular cameras [44, 45, 46, 49] are lightweight and cost-effective but restricted to 2D image data, relying on advanced algorithms for depth estimation. Stereo cameras utilize dual lenses to calculate depth through stereo triangulation, enabling 3D perception [50].
@@
RGB cameras, however, significantly degrade in performance under low-light or adverse weather conditions. To address these limitations, thermal and infrared cameras can serve as complementary sensors. Thermal cameras [^8] https://www.flir.com/oem/adas/adas-dataset-form/, which detect heat signatures, are highly effective for detecting pedestrians and animals in complete darkness or adverse weather conditions. However, their shorter range and lack of fine detail limit their utility for tasks like detecting road signs and require specialized deep learning solutions. Infrared cameras enhance visibility in low-light conditions and are suitable for detecting road signs, detecting and tracking pedestrians and vehicles [52]. Nonetheless, they are sensitive to bright light sources, which can cause glare and blooming artifacts. This issue can be mitigated with filters designed to reduce glare and enhance image quality in challenging scenarios. Despite their individual limitations, combining thermal or infrared cameras with RGB cameras creates a more robust sensor suite, ensuring reliable operation across diverse environments.
@@
Kim et al. [56] presented a model to classify agent types (pedestrian, cyclist, or vehicle) using radar data. Radar signals are transformed into range-velocity (RV) images that capture an object’s distance and motion. These RV images are processed by a convolutional LSTM to learn spatial and temporal features, followed by a fully connected layer to map these features to target classes. Steinhauser et al. [57] proposed a system focused on individual limb movements for pedestrian behavior prediction. Their method employs a deconvolution algorithm combined with a clustering technique to extract and analyze reflections from pedestrian body parts, decomposing the range-Doppler spectrum to identify the torso, arms, and legs. Radar typically complements LiDAR and camera inputs in SOTA datasets, providing reliable velocity and heading cues.
Rasterization. Rasterization converts raw sensor data into structured, grid-like representations, enabling prediction models to effectively process spatial and temporal information. For example, LiDAR point clouds are can be transformed into grid-based formats, such as intensity or height maps, that capture spatial geometry. Radar data adds dynamic attributes such as velocity and direction [58], which can be encoded either through color coding (e.g., higher speeds represented by brighter colors) [59] or as separate velocity channels, storing velocity values in an additional data layer for each grid cell. Similarly, camera images can be rasterized to map rich semantic information, such as object classes and bounding boxes. This unified rasterized format facilitates the fusion of multimodal sensor data.

### v3

@@ Abstract
As the potential for autonomous vehicles to be integrated on a large scale into modern traffic systems continues to grow, ensuring safe navigation in dynamic environments is crucial for smooth integration. To guarantee safety and prevent collisions, autonomous vehicles must be capable of predicting the future trajectories of surrounding traffic agents. Over the past decade, significant efforts from both academia and industry have been dedicated to designing solutions for precise trajectory forecasting. These efforts have produced a vast range of approaches, raising questions about how they differ, which challenges they address, and what limitations still remain in trajectory prediction. This paper reviews a substantial portion of recent trajectory prediction methods, proposing a taxonomy to classify existing solutions. It also provides a general overview of the prediction pipeline, covering input and output modalities, modeling features, performance evaluation metrics, and existing prediction paradigms. In addition, the paper discusses active research topics within the forecasting realm, addresses the posed research questions, and highlights the remaining research gaps and challenges.
@@ 1 Introduction
The era of autonomous driving is on the verge of revolutionizing the modern transportation system. Automotive companies like Tesla, Waymo, and Cruise are pioneers in today’s self-driving vehicle technology. The practical, innovative solutions developed by these companies, in conjunction with academic research, have led to rapid advancements in perception, prediction, planning, and control algorithms for autonomous driving. Predicting the future trajectories of surrounding traffic participants is essential for safe navigation and serves as a core task of the prediction module in the autonomous stack. The predicted trajectories are passed as input to planning algorithms to eliminate paths that may lead to potential collisions and to select a safe course of action. Due to the diversity of traffic scenarios, road layouts, interactions between traffic participants, various maneuvers, inherent randomness in human actions, and computational time requirements, the trajectory forecasting problem presents a range of challenges (refer to Figure 1 for factors impacting trajectory forecasting).
Autonomous vehicles (AVs) operate in heterogeneous traffic environments, and their prediction modules must account for the motion dynamics of all traffic participants, including vehicles, cyclists, pedestrians, and motorbikes. Among these, the motion of vehicles is influenced not only by their past motion, road topology, and physical constraints, but also by the behavior of nearby agents. As human drivers intuitively understand such interdependencies, e.g., anticipating that a neighboring car being cut off may respond with a sudden maneuver, AVs must be equipped with a similar interaction awareness mechanism. To equip AVs with this capability, the prediction module needs to build a representation of each target agent’s surroundings and account for the interdependencies among traffic participants’ decisions.
As for human motion, beyond the complexity of interaction and inter-dependency, predicting human trajectories involves accounting for complex human behavior. Pedestrian motion is influenced not only by past trajectory, road topology and the actions of surrounding agents but also by social norms, destination intent, and inherent behavioral randomness, e.g., stopping abruptly to check a phone, changing direction to greet someone, crossing at undesignated locations, or hesitating mid-crossing due to uncertainty or distraction. For precise pedestrian trajectory forecasting, impacting factors can be inferred from scene contextual information and general knowledge, e.g., a child and an adult exhibit different motion patterns. However, incorporating scene context increases computational demand, as it requires additional architecture blocks to extract relevant semantic and situational cues.
Additionally, prediction models that incorporate traffic rules can benefit from a reduced set of plausible future trajectories that comply with these constraints. However, human drivers may violate traffic rules by over-speeding, changing lanes over a solid line, turning from a non-designated lane, ignoring stop signs, or even running a red light. Returning to the aspect of human behavioral randomness, prediction models must account for the possibility of such violations to ensure robust and reliable performance in real-world deployment. Lastly, trajectory prediction is inherently multimodal, i.e., the same past trajectory can lead to multiple plausible futures. For example, a vehicle approaching an intersection in a lane marked for both straight and left-turn movements may proceed forward or make a left turn. Ideally, the prediction model should capture this multimodality and produce outputs in a format that can be easily integrated into downstream planning modules.
@@
Thus, the trajectory prediction task is shaped by a set of impacting factors that must be captured from the input data and represented in the model. To address these challenges, the research community has focused on designing prediction models capable of learning complex motion patterns and incorporating relevant contextual cues. Among the top-performing solutions is Trajectron++ [1], which integrates environmental context through a graph-structured recurrent architecture and Spatio-Temporal Transformer Network (S2TNet) [2], leveraging transformer layers to capture both spatial and temporal dependencies. Further, from the intersection between academia and industry, the MultiPath++ model [3], released by the Waymo research group, employs multi-context gating fusion and dynamically learns reference points (anchors) to support goal-oriented predictions. Collectively, these models combine diverse techniques, such as graph-based interaction modeling, transformer architectures, and dynamic goal sampling, leading to more accurate and robust trajectory predictions.
@@
On the other side, the commercial research prioritizes real-time execution, cost-effectiveness, handling uncertainty at all stages of the autonomous stack and generalization for robust operating. This leads each company to use its own combination of sensors, compatible with the set of algorithms it employs. Tesla’s prediction algorithm is an end-to-end vision-based deep learning model [^1] https://www.tesla.com/support/autopilot, forecasting trajectories based on the visual sequences and aggregated contextual information over time. The model continuously learns from the vast amount of data collected by Tesla’s global fleet of vehicles. In contrast, Waymo utilizes a combination of high-fidelity sensors (LiDAR, cameras, and radar). The sensor data is processed by the perception module and then passed to motion prediction and planning components. Waymo adopts a modular structure, ensuring that each component can be thoroughly tested independently and then integrated into a single pipeline [^2] https://waymo.com/research/. It is challenging to determine the specific algorithms employed in real operations, but Waymo has a strong R&D focus, actively publishing its latest research findings. Aurora company focuses on autonomous solutions for trucks serving logistics purposes. Its prediction module [^3] https://blog.aurora.tech/ performs joint object detection and trajectory forecasting, combining learning-based methods with rule-based constraints to align predictions with driving rules. To ensure safety across a wide range of scenarios, from highway driving delivering cargo between cities to urban logistics, Aurora’s system is tested extensively through a combination of virtual testing suites and real-world operations. Each company’s strategy reflects a balance between predictive accuracy and operational constraints, tailored to the specific purpose of their vehicles.
@@
On the testing side of autonomous driving algorithms, two robust open-source platforms provide an environment for evaluation: Autoware [^4] https://github.com/autowarefoundation/autoware and Apollo [^5] https://github.com/ApolloAuto/apollo. Autoware’s perception stack includes functionalities such as object recognition (2D/3D detection using LiDAR, cameras, and radar), obstacle segmentation (distinguishing drivable areas via point cloud processing), occupancy grid mapping (representing occupied and free spaces), and traffic light recognition (detecting light states and directions). However, it lacks an integrated prediction module. Starting from release 5.0, Apollo incorporates an integrated prediction module [^6] https://github.com/ApolloAuto/apollo/blob/master/modules/prediction/README.md, supporting a range of predictors: Free Move Predictor for non-lane-constrained obstacles, Single Lane Predictor for vehicles on a single lane, and Lane Sequence Predictor for multi-lane and behavior-based trajectory predictions. Later versions, such as 6.0.0 and 7.0.0, introduced advanced prediction models, including a semantic map-based pedestrian prediction model and the Inter-TNT prediction model with interactive prediction and planning evaluation. Thus, Apollo provides a comprehensive platform for testing prediction algorithms in the closed-loop setting — within the full autonomous driving stack. For iterative integration, a lightweight platform that couples perception and prediction is needed for rapid experimentation. It can expedite the development of error-handling mechanisms and help standardize the evaluation of prediction models under perception noise. Adding a joint learning feature would enable the prediction module to adapt to perception uncertainties while providing feedback to improve perceptual representations.
@@
To address the posed research questions, this study surveys the latest advancements in prediction models and offers a structured framework for understanding the trajectory forecasting landscape. We review key factors that influence the performance and robustness of prediction models. This analysis aims to assist researchers in identifying gaps in the current literature and guiding future developments in the field. Our contributions are as follows:
@@
    Provide an open repository [^7] https://github.com/AV-Lab/Trajectory_Prediction_Models_Zoo featuring two prediction pipelines to facilitate experimentation and benchmarking. The first pipeline includes multiple learning-based models that can be trained on SOTA or custom datasets and the second pipeline implements a detect–track–predict approach that supports several sensor configurations (e.g., frontal camera, stereo camera, LiDAR, LiDAR coupled with cameras).
@@ 1.1 Related Surveys
Figure 2: Structure of the review paper: the study comprises seven main sections. The first two sections - an overview of the prediction pipeline and a survey of modeling approaches, provide general knowledge of the trajectory prediction domain. Next two sections are dedicated to incorporating driving knowledge and addressing uncertainty focusing improving performance, robustness and generalization of prediction models. Last three sections cover planning-conditioned prediction, VLM-guided prediction, and collaboration between intelligent vehicles for trajectory prediction.
@@ 1.2 Scope and Structure
The scope of this review is defined by the objective to offer a comprehensive overview of the trajectory prediction field for novice researchers and provide discussion of active topics for already experienced practitioners. Figure 2 illustrates the structure of our review, detailing the sections and main points discussed.
The first part consists of two sections aimed at providing general knowledge of the trajectory prediction realm. Section 2 provides an overview of a trajectory prediction pipeline and covers problem formulation and prediction paradigms. Section 3 lays out the taxonomy and categorization of the modeling approaches and elaborates on the models available in the literature from each category. The next part is dedicated to the discussion on improving models’ performance, robustness and generalization. This part also comprises two sections. Section 4 reviews the inclusion of driving knowledge in prediction models, covering different types of driving knowledge and the ways they are incorporated into prediction models. Section 5 is dedicated to uncertainty estimation and mitigation in prediction models. It elaborates on the sources of uncertainty and approaches to mitigate the risks associated with them. The last part depicts active research areas and is presented by three sections. Section 6 discusses planning-prediction-coupled frameworks, where for generating predictions the ego-vehicle models internal planning intentions of surrounding agents. Next, Section 7 discusses the application of Vision-Language models in scene understanding and trajectory prediction. Further, Section 8 discusses the concept of collaborative perception, elaborating on different types of fusion and different types of challenges caused by collaboration. Lastly, Section 9 provides a discussion on the posed research questions, outlines the research gaps, and highlights future research directions.
@@ 2.1 Problem Formulation
In the given formulation, the predictor ($\mathcal{P}$) is a function of a single parameter, past trajectories ($\mathcal{T}$). This can be expanded to make $\mathcal{P}$ a function of multiple parameters if additional inputs are available, e.g., road geometrical features ($G$) or interaction modeling features ($I$). Additionally, the coordinate component of $\tau$ can be represented in either a local or global coordinate system. Sensor observations are inherently centered on the ego-vehicle, meaning all raw measurements are initially recorded in the vehicle’s local frame. As such, predicting trajectories in the local frame does not require coordinate transformations. However, local frames introduce ego-motion distortions like stationary vehicles may appear to move, and turning maneuvers can be misrepresented due to the ego-vehicle’s own motion. To ensure consistent learning of motion patterns, ego-motion must be compensated, e.g., through frame-to-frame undistortion using odometry interpolation or local frame stabilization using ego-motion data. Alternatively, trajectories can be transformed to a global frame using ego-pose information, as done in [17], though this requires precise localization.
Representing trajectories in Cartesian coordinates ties them to the underlying road geometry, making it more challenging for prediction models to generalize across scenes where structurally different layouts exhibit similar behaviors. The Frenet coordinate frame addresses this by encoding each trajectory point relative to the ego-vehicle’s path as $(s,d)$, where $s$ denotes the arc length along the lane centerline and $d$ the perpendicular lateral offset. This decouples motion patterns from the global layout and enables learning behaviors relative to the road structure [18].
@@ 2.2.1 Past Trajectories
Predicting future trajectories solely based on past trajectories [19, 20, 21, 22] forces the models to learn patterns by relying purely on spatio-temporal dependencies. Predicting future trajectories solely based on past motion is efficient in terms of prediction speed and computational cost. One way to incorporate additional cues is to implicitly encode interaction of the target agent. This can be done by including the past trajectories of its surrounding agents in the prediction process, as shown in the studies [23, 24, 25, 26, 27, 28, 29]. This enables predictions that better reflect real-world behaviors, such as reactions to nearby vehicles or pedestrians. However, when forecasting models are trained on a single input source under the assumption of error-free observations, their performance will degrade in real-world deployments. In practice, the past trajectories available to prediction models are often noisy and incomplete. Therefore, models should be evaluated under such conditions (e.g., by simulating tracker errors) to quantify the magnitude of prediction error, assess sensitivity, and evaluate the model’s robustness to inaccurate inputs. On representations, typically the agent state in past trajectories is represented as $(x,y)$ coordinates and models are trained on relative distances. Additionally, heading and speed can be added complementarily. Studies [30, 31, 32] improved prediction accuracy by including velocity, acceleration, and orientation as input features.
@@ 2.2.2 Sensors Data
Sensor data from LiDAR, radar, or cameras is used as prime or additional input to prediction models. This section first examines each sensor individually and then discusses approaches of integrating multi-modal inputs.
@@
LiDAR. LiDAR sensors enable environmental perception through laser-based distance measurements, making them indispensable for autonomous driving systems. Two types of LiDAR are commonly employed: perimeter LiDAR and $360^{\circ}$ LiDAR. Perimeter LiDAR, mounted on the vehicle’s edges or corners, is optimized for monitoring close-range objects, blind spots, and collision risks, particularly in scenarios like parking or navigating tight spaces. In contrast, $360^{\circ}$ LiDAR, typically roof-mounted, provides a complete $360^{\circ}$ field of view, supporting long-range detection and situational awareness. Current SOTA datasets predominantly include data from $360^{\circ}$ LiDARs (refer to Figure 4 for examples), while perimeter LiDARs are used by automotive manufacturers as complementary sensors to enhance overall perception. One of LiDAR’s key strengths is its ability to function reliably in all lighting conditions, including complete darkness and bright sunlight, due to its light-invariant nature. However, its performance degrades in adverse weather, such as heavy rain, or snow, where laser beam scattering reduces detection range and accuracy along with highly reflective surfaces such as glass, potentially creating blind spots. Despite these challenges, LiDAR remains highly effective for detecting obstacles [33], vehicles [34, 35, 36], lane markers [37, 38, 39, 40], and forecasting [41].
@@
Two limitations hinder the widespread adoption of LiDAR technology: its relatively high cost and its accuracy in detecting pedestrians. Although LiDAR technology is advancing and prices are gradually declining, it remains expensive and impractical for consumer-level mass-market adoption. The second limitation is its lower accuracy in recognizing fine-grained classes such as pedestrians and cyclists. This disparity stems from its sparse, geometry-centric output and lack of rich semantic information, which makes interpretation of complex objects challenging. For example, on the KITTI leaderboard, the top pedestrian detection score is only 56% [44], compared to a 93% vehicle detection score achieved by [45]. As a result, LiDAR systems are ideally complemented by cameras, which provide the necessary semantic context to enhance pedestrian detection performance.
Camera. Cameras play a key role in perception of autonomous vehicles in detecting traffic participants, road signs, and traffic lights. These sensors are characterized by their resolution, field of view, and configuration. The vision community is highly advanced, providing efficient algorithms for object detection [46], semantic segmentation [47], and lane recognition [48]. Camera setups in AV systems are generally categorized as single frontal cameras [49] or multi-view configurations [50]. Single frontal camera setups primarily rely on 2D perception and are mostly used in driving assistance systems. In contrast, multi-view setups [51] arrange cameras to ensure complementary, non-overlapping fields of view, providing 360-degree observability. Additionally, camera systems employ either monocular or stereo setups. Monocular cameras [46, 47, 48, 51] are lightweight and cost-effective but restricted to 2D image data, relying on advanced algorithms for depth estimation. Stereo cameras utilize dual lenses to calculate depth through stereo triangulation, enabling 3D perception [52].
@@
RGB cameras, however, degrade in performance under low-light or adverse weather conditions. To address these limitations, thermal and infrared cameras can serve as complementary sensors. Thermal cameras [^8] https://www.flir.com/oem/adas/adas-dataset-form/, which detect heat signatures and can be utilized in detecting pedestrians and animals in complete darkness or adverse weather conditions. However, their shorter range and lack of fine detail limit their utility for tasks like detecting road signs and require specialized deep learning solutions. Infrared cameras enhance visibility in low-light conditions and are suitable for detecting road signs, detecting and tracking pedestrians and vehicles [54]. Nonetheless, they are sensitive to bright light sources, which can cause glare and blooming artifacts. This issue can be mitigated with filters designed to reduce glare and enhance image quality in challenging scenarios. Despite their individual limitations, combining thermal or infrared cameras with RGB cameras creates a more robust sensor suite, ensuring reliable operation across diverse environments.
@@
Kim et al. [58] presented a model to classify agent types (pedestrian, cyclist, or vehicle) using radar data. Radar signals are transformed into range-velocity (RV) images that capture an object’s distance and motion. These RV images are processed by a convolutional LSTM to learn spatial and temporal features, followed by a fully connected layer to map these features to target classes. Steinhauser et al. [59] proposed a system focused on individual limb movements for pedestrian behavior prediction. Their method employs a deconvolution algorithm combined with a clustering technique to extract and analyze reflections from pedestrian body parts, decomposing the range-Doppler spectrum to identify the torso, arms, and legs. Radar typically complements LiDAR and camera inputs in SOTA datasets, providing velocity and heading cues.
Rasterization. Rasterization converts raw sensor data into structured, grid-like representations. For example, LiDAR point clouds can be transformed into grid-based formats, such as intensity or height maps, that capture spatial geometry. Radar data adds dynamic attributes such as velocity and direction [60], which can be encoded either through color coding (e.g., higher speeds represented by brighter colors) [61] or as separate velocity channels, storing velocity values in an additional data layer for each grid cell. Similarly, camera images can be rasterized to map rich semantic information, such as object classes and bounding boxes. This unified rasterized format facilitates the fusion of multimodal sensor data.